OpenStudy (anonymous):

Help me find the domain of this logarithmic f(x)=log[3](x-2)

OpenStudy (***[isuru]***):

the log values r not defined for the negative numbers... so x - 2 > 0 x > 2

OpenStudy (***[isuru]***):

so... the domain will be.. \[[ +\infty > x > 2 \ ]\]

OpenStudy (***[isuru]***):

got it ?

OpenStudy (anonymous):

so the answer would be \[2,\infty\]

OpenStudy (***[isuru]***):

yeah but u have to be concern about the brackets.... they should be () not []

OpenStudy (anonymous):

okay thanks alot :D

OpenStudy (***[isuru]***):

srry.. i made a mistake it's \[( 2 , + \infty )\]

OpenStudy (anonymous):

x-2<0 x>2 is this it?

OpenStudy (***[isuru]***):

yeah

OpenStudy (anonymous):

okay :D thanks again

OpenStudy (***[isuru]***):

u r welcome!